OpenStudy (anonymous):

Solve. 3/x = 1/x-2 A. x=3 B. x=4 C. x=6

OpenStudy (anonymous):

3 works by inspection

OpenStudy (anonymous):

\[\frac{3}{3}=\frac{1}{3-2}\] \[1=1\]

OpenStudy (anonymous):

answer is A